# Node-like as a Whole: Structure-aware Searching and Coarsening for Graph Classification

### v2

@@ 5.4.1. Time Analysis
Figure 6. Runtimes of all coarsening algorithms. We make logarithmic processing due to different scales of datasets.
@@
Table 5. Scale analysis of graph coarsening. Each group represents the size of the intermediate coarsening graph under different strategies, where $\bar{V}$ and $\bar{E}$ represent the average nodes and edges, and $r_{V}$ and $r_{E}$ represent the scaling ratio of the coarsening graph compared to the original graph. The negative number means reducing the scale from the original graph, and the positive means expanding.
Method CO IB IM DD N1 PTC PRO N109
Original $\bar{V}$ 74.49 19.77 13.00 284.32 39.06 25.56 29.87 29.68
$\bar{E}$ 2457.50 96.53 65.94 715.66 72.82 25.96 32.30 32.13
nx.cycle $\bar{V}$ 74.49 19.77 13.00 284.32 29.84 25.30 39.06 29.65
@@
We analyze the linear time complexity of LCC in section 4.2.4. Here, we compare it with other methods. Figure 6 shows the results.
The results show that the runtime of LCC is significantly lower than other methods. Except for CO and DD, the computation time is within 2s for all other datasets. Runtime for CO is 7.60s, while for DD is 37.58s. The average runtime of KGC.neighbor is 44.55s, and L19.clique is 54.98s. Both of them are higher than LCC.
5.4.2. Scale Analysis
@@
(a) Loops: N109 (77)
(b) Cliques: IB (14)
(c) Plain: PTC (5)
Figure 7. Visualization of different coarsening strategies. For (a) loops, (b) cliques, and (c) plain graphs without loops and cliques, we find corresponding sample graphs from datasets (graph ID in parentheses). In each group, row 1 col 1 is the original input graph, row 2 col 1 is the coarsening result of GRLsc, and the other columns are two variants of the method in the same category. Different node colors represent different node classes. The graph layout does not contain semantics such as rotation and symmetry but only shows the connections.
In general, the LCC we designed can achieve a considerable coarsening result. The coarsening rate at the node level is about 52.2%. It is slightly higher at the edge level, 57.9% or so. In baselines, the best one is KGC.neighbor, with a 70.5% compression rate at the node level and 54.7% at the edge level. We do not expect LCC to achieve similar high rates. When the coarsening process is to the maximum extent, amounts of high-level structural information are lost.
Table 6. Optimal parameter settings.
Hyper-parameters CO IB IM DD N1 PTC PRO N109
$K$ 1 1 3 2 1 3 2 1
$T$ 2 2 1 2 2 4 1 1
$N$ 16 16 4 16 16 16 4 8
$|H|$ 1024 1024 1024 1024 1024 1024 1024 1024
$Lr$ 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
$|B|$ 4 4 4 4 4 4 4 4
Figure 8. Hyper-parameters Study. We evaluate four hyper-parameters: hierarchy depth $K$, the number of Transformer Encoder layers $T$, the number of sampling neighbors $N$, and the hidden size $|H|$.
As the number of nodes decreases linearly, edges have a faster decay. We cannot mine enough structural information in very simple hypergraphs. Thus, LCC achieves the maximum balance between coarsening nodes and preserving structures.
5.4.3. Visualization
We further analyze the coarsening comparison using visualization. Figure 7 shows the results.
Figures 7(a) and  7(b) visualize the loops and cliques that GRLsc focuses on. It appears that GRLsc can obtain the coarsening graph representation closest to the original graph. We retain the connections between supernodes to the greatest extent and mine the high-level structural information thoroughly based on the coarsening procedures. Other algorithms lack balance in coarsening and structural representation. Random algorithms break away from the original structural semantics and pursue to cover the whole graph by specifying the number of nodes. KGC presents a more concise coarsening pattern. However, it fails to express high-level structural information and has poor adaptability to loops. L19 relaxes the restriction, not limiting nodes located in multiple partitioning sets, but still cannot refine the structures.
In addition, we analyze for plain graphs that do not contain loops and cliques, as shown in Figure 7(c). GRLsc returns the same graph $G_{lcc}$ as the original input $G_{o}$ through LCC and makes a further conversion via LGC. In this kind of graph, KGC can effectively identify other structures as claw-like ones and realize their coarsening process.
5.5. Hyper-parameters Study
As for $K$ and $T$, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy. The most striking example is the plot of CO and DD on $T$, where the performance of GRLsc decreases substantially when the number of layers changes from 3 to 4.
5.6. Case Study
@@
Figure 9. Case study for all eight datasets. Each case contains three graphs from left to right: the original graph view, the graph coarsening view, and the line graph conversion view. The heat plot below them is the classifier weights, with the x-axis representing the feature dimension and the y-axis representing the three graphs (each graph view takes rows of the number of graph classes).
We conduct case studies on all eight datasets to fully cover each category and explore the choices of the classifier after GRLsc. The results are in Figure 9.
It can be seen from the cases intuitively that different datasets have different structural characteristics. Even datasets in the same category also have differences in representation. For example, (a) CO, (c) IB, and (d) IM have strong clique structure distribution, graphs in (h) PRO contain multiple connected branches and the similar datasets (e) N1 and (f) N109 consist of multi-chain and multi-loop structure, respectively, and so on. GRLsc focuses on loops and cliques and thus performs well on datasets rich in such structures. As for loops, we take (f) N109 as an example, which contains three benzene rings connected in turn. LCC can identify such loop structures and coarsen them into three supernodes. The remaining two independent carbon atoms are then converted by LGC into a triangle, as shown in the figure. We can see from the heatmap that both LCC and LGC contain vital dimensions contributing most to the classification weights, where LCC in (35, 3), LGC in (6, 4), (24, 5), etc.
@@
In addition to examples with remarkable features, the dataset also contains (b) DD with a more complex structure and (e) N1 without loops and cliques. We can observe the unique role of LCC and LGC from them. The former coarsens the loops and cliques via LCC and significantly reduces the graph scale, while the latter shortens long chains and marks the position of special forks, such as claw-like structures, through LGC.
Across all eight datasets, we notice that (h) PRO has more unique properties. The original graph is possibly not connected, containing some independent connected branches or components. LCC and LGC start from one particular node, traversing the node and edge set completely. Therefore, multiple connected branches do not affect the operation of our algorithm.
@@
6. Conclusion and Future Work
6.1. Conclusion
In this paper, we propose a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc) on GT architectures for graph classification task. We focus on loops and cliques and investigate the feasibility of treating particular structures node-like as a whole. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of GRLsc on six real-world datasets, and the experimental results demonstrate that GRLsc can outperform SOTAs in graph classification task.
6.2. Future Work
Though GRLsc achieves remarkable results, we still have a long way to go. In the future, there are two main directions for discussion. First, GRLsc prefers to deal with static graphs at present, while graphs in the real world are constantly changing over time. We try to introduce dynamic graphs. Second, GRLsc focuses on two specific structures: loops and cliques. However, graph structures are more complex and diverse. We will consider extending to general structures to mine richer information at a high level.

### v5

@@ 5.4 Hyper-parameters Study
Figure 5: Runtimes of all coarsening algorithms. We make logarithmic processing due to different scales of datasets.
5.5 Coarsening Strategy Ablation
To better discuss the impact of different coarsening strategies on the model performance, we select four categories of graph coarsening schemes: random, networkx (N) [65], KGC (K) [28], and L19 (L) [30]. Under each category, there are two variants: neighbor (*n) and clique (*c).
In terms of implementation, we set $K=5$ and $r=0.5$, according to the paper [28, 30]. The algorithm strategy is greedy, and other parameters are consistent with the origin. To ensure alignment with LCC4GC, the maximum coarsening level is set to 1, which means only coarsening the original graph without higher-level compressions. The random algorithm sets $R=K=5$, and each supernode randomly selects components from the sampling sets until all nodes are covered. Table 6 shows the model performance under different coarsening strategies, where LCC achieves the optimal results under all eight datasets. It indicates that the graph coarsening view obtained by LCC is more suitable for graph classification tasks.
5.5.1 Time Analysis
Table 6: Coarsening strategy ablation studies on all eight datasets. LC is the abbreviation of LCC.
Method CO DD PTC PRO
Neighbor $98.12^{0.65}$ $98.64^{1.94}$ $70.67^{6.53}$ $75.20^{3.99}$
Random $97.94^{0.58}$ $99.32^{0.74}$ $71.81^{4.61}$ $72.51^{5.68}$
Nn $98.02^{0.60}$ $86.16^{4.10}$ $75.92^{6.31}$ $74.30^{4.57}$
Nc $98.04^{0.58}$ $86.19^{5.62}$ $72.05^{6.67}$ $72.77^{6.16}$
Kn $98.08^{0.58}$ $97.71^{2.57}$ $75.00^{2.66}$ $73.13^{3.70}$
Kc $96.06^{0.77}$ $82.00^{7.04}$ $70.61^{8.86}$ $74.75^{5.05}$
Ln $96.92^{0.84}$ $96.44^{4.82}$ $74.06^{8.07}$ $74.21^{4.45}$
Lc $97.28^{0.66}$ $91.08^{3.00}$ $71.20^{6.10}$ $74.30^{3.65}$
LC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\textbf{76.37}}^{4.37}$
Method IB IM N1 N109
Neighbor $89.10^{4.09}$ $62.80^{3.45}$ $77.79^{1.48}$ $76.40^{1.99}$
Random $89.10^{3.78}$ $61.70^{1.90}$ $77.79^{2.36}$ $74.10^{1.80}$
Nn $87.00^{2.31}$ $62.47^{4.06}$ $74.84^{2.59}$ $73.81^{2.21}$
Nc $87.90^{2.62}$ $62.53^{3.41}$ $75.69^{1.92}$ $73.88^{1.81}$
Kn $87.00^{1.95}$ $58.87^{4.03}$ $76.18^{2.12}$ $73.42^{1.36}$
Kc $89.30^{2.72}$ $59.40^{3.66}$ $74.60^{1.91}$ $72.84^{1.72}$
Ln $87.10^{2.62}$ $61.70^{3.17}$ $75.94^{2.20}$ $73.81^{2.14}$
Lc $88.20^{2.79}$ $61.87^{3.55}$ $77.81^{1.48}$ $75.14^{1.86}$
LC $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\textbf{79.08}}^{1.92}$ $\underline{\textbf{76.98}}^{1.90}$
We analyze the linear time complexity of LCC in section 4.2.4. Here, we compare it with other methods. Figure 5 shows the results. The results show that the runtime of LCC is significantly lower than other methods. The runtime of LCC is significantly lower than other methods. Except for CO and DD, the computation time is within 2s for all other datasets. Runtime for CO is 7.60s, while for DD is 37.58s. The average runtime of KGC.neighbor is 44.55s, and L19.clique is 54.98s. Both of them are higher than LCC.
@@ 5.5.1 Time Analysis
Table 7: Scale analysis of graph coarsening. Each group represents the size of the intermediate coarsening graph under different strategies, where $\bar{V}$ and $\bar{E}$ represent the average nodes and edges, and $r_{V}$ and $r_{E}$ represent the scaling ratio of the coarsening graph compared to the original graph. The negative number means reducing the scale from the original graph, and the positive means expanding. O is the original graphs.
CO IB IM DD N1 PTC PRO N109
O $\bar{V}$ 74.49 19.77 13.00 2.8e2 39.06 25.56 29.87 29.68
$\bar{E}$ 2.5e3 96.53 65.94 7.2e2 72.82 25.96 32.30 32.13
Nn $\bar{V}$ 74.49 19.77 13.00 2.8e2 29.84 25.30 39.06 29.65
@@
5.5.2 Scale Analysis
@@
In general, the LCC we designed can achieve a considerable coarsening result. The coarsening rate at the node level is about 52.2%. It is slightly higher at the edge level, 57.9% or so. In baselines, the best one is KGC.neighbor, with a 70.5% compression rate at the node level and 54.7% at the edge level. We do not expect LCC to achieve similar high rates. When the coarsening process is to the maximum extent, amounts of high-level structural information are lost. As the number of nodes decreases linearly, edges have a faster decay. We cannot mine enough structural information in very simple hypergraphs. Thus, LCC achieves the maximum balance between coarsening nodes and preserving structures.
5.5.3 Visualization
(a) Cliques: IB (14)
(b) Loops: N109 (77)
Figure 6: Visualization of different coarsening strategies. In each group, row 1 col 1 is the original input graph, row 2 col 1 is the coarsening result of LCC4GC, and the other columns are two variants of the method in the same category. Different node colors represent different node classes. The graph layout does not contain semantics such as rotation and symmetry but only shows the connections.
We further analyze the coarsening comparison using visualization. Figure 6 shows the results. It appears that LCC4GC can obtain the coarsening graph representation closest to the original graph. We retain the connections between supernodes to the greatest extent and mine the high-level structural information thoroughly based on the coarsening procedures. Other algorithms lack balance in coarsening and structural representation. Random algorithms break away from the original structural semantics and pursue to cover the whole graph by specifying the number of nodes. KGC presents a more concise coarsening pattern. However, it fails to express high-level structural information and has poor adaptability to loops. L19 relaxes the restriction, not limiting nodes located in multiple partitioning sets, but still cannot refine the structures.
5.6 Case Study
@@
Figure 7: Case study for all datasets. Each case contains three graphs from left to right: the original graph view, the graph coarsening view, and the line graph conversion view. The heat plot below them is the classifier weights, with the x-axis representing the feature dimension and the y-axis representing the three graphs (each graph view takes rows of the number of graph classes).
We conduct case studies on all eight datasets to fully cover each category and explore the choices of the classifier after LCC4GC. Figure 7 illustrates the results.
LCC4GC focuses on loops and cliques and thus performs well on datasets rich in such structures. As for loops, we take (f) N109 as an example, which contains three benzene rings connected in turn. LCC can identify such loop structures and coarsen them into three supernodes. The remaining two independent carbon atoms are then converted by LGC into a triangle, as shown in the figure. We can see from the heatmap that both LCC and LGC contain vital dimensions contributing most to the classification weights, where LCC in (35, 3), LGC in (6, 4), (24, 5), etc.
@@
6 Conclusion
In this paper, we propose an efficient Loop and Clique Coarsening algorithm with linear complexity for Graph Classification (LCC4GC) on GT architectures. We focus on loops and cliques and investigate the feasibility of coarsening particular structures into hypernodes. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of LCC4GC on eight real-world datasets, and the experimental results demonstrate that LCC4GC can outperform SOTAs in graph classification task. Though LCC4GC achieves remarkable results, we still have a long way to go. In the future, there are two main directions for discussion. First, graphs in the real world are constantly changing over time, we will try to introduce dynamic graphs. Second, graph structures are more complex and diverse than just loops and cliques, we will consider extending to general structures to mine richer information at a high level.